OpenStudy (anonymous):

I WILL FAN AND MEDAL!!!!!!!!!! PLZ HELP!!!!!! If y varies inversely as x and the constant of variation is thirty, what is y when x equals six? Choose one answer. A. 5 B. twenty-four C. thirty-six D. one hundred eighty

OpenStudy (anonymous):

@Pawanyadav @MrNood

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (anonymous):

@lexidavi

OpenStudy (anonymous):

@jornelso @KyanTheDoodle @h@BlossomCake @Luigi0210 @cqluv @mathmath333

OpenStudy (anonymous):

@cramos725

OpenStudy (pawanyadav):

y=30/x So put x=6 I hope you understand why y=30/x

OpenStudy (anonymous):

30/6=5

OpenStudy (anonymous):

so is the answer A?

OpenStudy (pawanyadav):

Yes

OpenStudy (anonymous):

ok cool